OpenStudy (anonymous):

if they give me amplitude:3 phase shift: -1/6 and period: 5pi whats the equation for sine please help!

OpenStudy (phi):

\[Asin(\omega t+\phi)\] Also, angular frequency is \[\omega=2\pi f \] Frequency in Hertz is \[ f=\frac{1}{T}\] where T is the period. so: \[A= 3, f=\frac{1}{5 \pi}, \omega=2 \pi f= \frac{2}{5},\phi=-\frac{1}{6}\] \[3\sin(\frac{2}{5}t-\frac{1}{6})\]